# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ 3 Analysis
To proceed with our analysis, we first present some preliminary concepts and results. These are followed by a formal statement of the main results, then a discussion of their intuitions, proofs, and policy implications.
@@ 3.1 Preliminaries
We first formalize the problem faced by the DM. Let $\mathcal{N}(b,\Sigma)$ denote an arbitrary $N$-dimensional normal distribution with mean $b\in\mathbb{R}^{N}$ and covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$, and let $\mathcal{N}(b^{*},\Sigma^{*})$ denote the true distribution of the signals. The Kullback-Leibler divergence from $\mathcal{N}(b,\Sigma)$ to $\mathcal{N}(b^{*},\Sigma^{*})$ is
@@
The next lemma — which adapts the seminal result of Berk (1966) and that of HKS to the current setting — shows that the DM’s long-run belief concentrates on the set of distributions that minimize the Kullback-Leibler divergence from the true distribution (hereafter KL minimizers). That is, for every open cover of the set of KL minimizers, the DM assigns probability one to it in the limit, as $t\rightarrow+\infty$.
@@
[…]
Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ be a typical KL minimizer. Since $\Sigma^{*}$ is known to the DM from the outset, $\widehat{\Sigma}=\Sigma^{*}$. Simplifying the KL divergence accordingly turns the DM’s problem into:
@@
In Section 3.3, we show that the solution to (2) is unique and fully captured by $\Delta\coloneqq\hat{b}-b^{*}$. Call $\Delta$ the DM’s misperception about source biases in the long run, and note that $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}$ $\forall i\in\mathcal{M}$. Let $\text{BLP}^{*}(X)$ and $\widehat{\text{BLP}}(X)$ denote the best linear predictors of the state given signals, under the belief that source biases have values $b^{*}$ and $\hat{b}$, respectively. These are the long-run behaviors of the DM under the correctly specified model and the misspecified model, respectively, and their difference captures the distortionary effect of misspecified learning on the DM’s long-run behavior.
3.2 Results
Our main results are twofold. The first characterizes the distortionary effects of misspecified learning on the DM’s long-run belief and behavior. The second investigates comparative statics of distortionary effects.
For starters, write $\nu_{i}$ for the signal-to-noise ratio $1/v_{i}$ of source $i$. For each $i\in\mathcal{M}$, define
$\gamma_{i}\coloneqq\frac{\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$
as the relative precision of the signal generated by source $i$ compared to other misspecified sources. Define
$\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\gamma_{i}\Delta_{i}$
as the total misspecification across the sources in $\mathcal{M}$, weighted by their relative precision to each other. Then
@@ 3.2 Results
  1. (i)
    $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$; $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
  2. (ii)
    $\overline{\Delta}$ depends only on $(\Delta_{i},\nu_{i})_{i\in\mathcal{M}}$. For each $i\in\mathcal{M}$, $\partial\overline{\Delta}/\partial\Delta_{i}>0$, while $\partial\overline{\Delta}/\partial\nu_{i}$ generally has an ambiguous sign.
Part (i) of Theorem 1 reduces the distortionary effects of misspecified learning to a single metric: $\overline{\Delta}$. To develop intuitions, suppose that the manager in our leading example begins with an erroneous belief that two sources are both impartial, although in reality one is biased against minorities and the other is in favor of them, that is, $b_{1}^{*}<0<b_{2}^{*}$ and $\tilde{b}_{1}=\tilde{b}_{2}=0$. Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences. One potential approach is to find a middle ground, that is, a position in $(b_{1}^{*},b_{2}^{*})$, and use it as a neutral point to form beliefs about other sources. The more precise source $i$ is relative to source $j$, the closer the neutral point is to $b_{i}^{*}$. In the case where $\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}<0$, the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities of magnitude $\overline{\Delta}=|\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}|$, and it is mirrored in his beliefs about unfamiliar sources through misspecified learning. In the long run, the DM underestimates the bias of any source that is adversarial toward minorities while overestimating that of any source that is favorable to them, assuming that $\overline{\Delta}$ is not excessively large. He also underestimates minority employees’ productivity and pays them less than under the correctly specified model by $\overline{\Delta}$.
Part (ii) of Theorem 1 investigates the comparative statics of our metric. The result $\partial\overline{\Delta}/\partial\Delta_{i}>0$ is the most intuitive. In the leading example, this means that correcting the unconscious bias associated with any familiar source reduces its impact on the metric.
The comparative statics regarding source precision are more nuanced. As we increase the absolute precision of source $i$ (as captured by $\nu_{i}$), we also increase its relative precision (as measured by $\gamma_{i}$), while decreasing the relative precision $\gamma_{j}$ of any other source $j\in\mathcal{M}-\{i\}$. As a consequence, the DM takes source $i$ more seriously in the learning process, and misspecifications of its bias distort the DM’s long-run belief and behavior more heavily than before. The opposite happens to source $j$. Depending on the exact misspecifications associated with each individual source, the outcome is either an exacerbation of the distortionary effect, or a mitigation thereof. Such an ambiguity hinges on the multi-sourced nature of misspecifications, i.e., $M\geq 2$. If instead $M=1$ (as in HKS), then raising $\nu_{i}$ unambiguously increases the magnitude of $\overline{\Delta}$, due to the first effect described above.
Perhaps the most striking feature of $\overline{\Delta}$ is that it depends only on misspecified sources (DOOM) but not those with initially unknown biases. In general, DOOM fails because the correlations between the signals affect how the latter are weighed in the learning process. The presence of multi-sourced misspecifications can significantly complicate the analysis, as the need to weigh different sources promotes the role of correlation in shaping learning outcomes. The compound effect may escape even the most sophisticated analysis. When correlations are concentrated on only a few channels (as in the baseline model), their impacts on learning outcomes are limited. The proof presented in the next section formalizes these ideas.
3.3 Proof of Theorem 1
Throughout this paper, we use $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$, $\text{diag}(\cdot)$ to denote the square matrix created from a vector, and $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, for each $k\in\mathbb{N}$. Write the covariance matrix $\Sigma^{*}$ of the signals as
$\left[\begin{array}[]{c|c}A_{M\times M}&B_{M\times(N-M)}\\ \hline\cr C_{(N-M)\times M}&D_{(N-M)\times(N-M)}\end{array}\right],$
where
capture the correlations between familiar sources and those between familiar and unfamiliar sources, respectively.
By Lemma 1, the DM’s long-run belief is a solution to (2). Rewrite (2) as
@@ 3.3 Proof of Theorem 1
$\sum_{i=1}^{N}\Sigma^{*-1}_{ji}\Delta_{i}=0.$ (3)
We proceed in two steps.
Step 1.
Consider first the case where $\Sigma^{*}$ is an arbitrary symmetric positive definite matrix. The next lemma solves for (3) in this general case. For ease of notation, write $\widetilde{\Delta}$ for the vector $[\Delta_{1},\cdots,\Delta_{M}]^{\top}$ of misspecifications associated with familiar sources, and $\widehat{\Delta}$ for the vector $[\Delta_{M+1},\cdots,\Delta_{N}]^{\top}$ of misperceptions about unfamiliar sources.
Lemma 2.
For any symmetric positive definite matrix $\Sigma^{*}\in\mathbb{R}^{N\times N}$, the solution to (3) is $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ and in particular satisfies $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$.
The proof of Lemma 2 exploits only the first-order conditions associated with KL minimization, together with the block matrix inversion formula. The result prescribes a two-step procedure for aggregating the misspecifications associated with familiar sources into distortions of the DM’s long-run belief and behavior. The first step weighs misspecifications using the correlations between familiar sources (as captured by $A^{-1}\widetilde{\Delta}$). The second step transforms these weighted sums into the DM’s misperceptions about unfamiliar sources, using the correlations between familiar and unfamiliar sources as captured by $C$. Interestingly, correlations between unfamiliar sources do not enter this weighting process. Furthermore, the distortion of the DM’s long-run behavior depends only on misspecified sources but not those he learns about, thus always satisfying DOOM.
Step 2.
While the DM’s misperceptions about source biases do not generally satisfy DOOM due to the complications that arise in the weighting process (as captured by $CA^{-1}$), they become relatively straightforward when the correlations between the signals are concentrated on only a few channels. In the environment considered in the baseline model, distortions of the DM’s long-run perceptions and behavior collapse into a single metric: $\overline{\Delta}$.
A seemingly innocuous assumption turns out to be key, namely signals net of their biases equal the true state plus idiocyncratic errors. As a result, their covariance matrix $\Sigma^{*}$ is the sum of a diagonal matrix: $\text{diag}(v)$, and a matrix of rank one: $\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. The next lemma, proven by Sherman and Morrison (1950), prescribes a formula to invert rank-one updates of invertible matrices.
Lemma 3 (Sherman and Morrison 1950).

### v3

@@ 3 Analysis
To proceed with our analysis, we first present some preliminary concepts and results. These are followed by a formal statement of the main results, then a discussion of their intuitions, proofs, and implications.
@@ 3.1 Preliminaries
Throughout the paper, we follow the convention of using $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$. For each $k\in\mathbb{N}$, we use $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, and $\mathbb{I}_{k\times k}$ to denote the $k\times k$ identity matrix.
We first formalize the DM’s problem. Let $\mathcal{N}(b,\Sigma)$ denote an arbitrary $N$-dimensional normal distribution with mean $b\in\mathbb{R}^{N}$ and covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$, and let $\mathcal{N}(b^{*},\Sigma^{*})$ denote the true distribution of the signals. The Kullback-Leibler divergence from $\mathcal{N}(b,\Sigma)$ to $\mathcal{N}(b^{*},\Sigma^{*})$ is
@@
The next lemma adapts the seminal result of Berk (1966) and, more recently, Theorem A.1 Case (I) of HKS, to the current setting, showing that the DM’s long-run beliefs concentrate on the set of distributions that minimize the Kullback-Leibler divergence from the true distribution (hereinafter, KL minimizers). That is, for every open cover of the set of KL minimizers, the DM assigns probability one to it in the limit, as $t\rightarrow+\infty$.
@@
The DM’s long-run beliefs concentrate on the set of distributions that minimize the KL divergence (1), subject to the constraint that $b\in\Theta$ and $\Sigma=\Sigma^{*}$.
Proof.
Omitted proofs from the main text are relegated to the appendix. ∎
Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ denote a typical KL minimizer. Since $\Sigma^{*}$ is known to the DM from the outset, $\widehat{\Sigma}=\Sigma^{*}$. Simplifying the KL divergence accordingly turns the DM’s problem into:
@@
In the next section, we show that the solution to (2) is unique and fully captured by $\Delta\coloneqq\hat{b}-b^{*}$. Call $\Delta$ the DM’s misperception about source biases in the long run, and note that $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\neq 0$ for all $i\in\mathcal{M}$. Let $\text{BLP}^{*}(X)$ and $\widehat{\text{BLP}}(X)$ denote the best linear predictors of the state given signals, under the beliefs that source biases have values $b^{*}$ and $\hat{b}$, respectively. These are the long-run predictions made by the DM under the correctly specified and misspecified models, and their difference captures the distortionary effect of misspecified learning on the long-run prediction.
In the leading example, $\Delta$ represents the DM’s misperception about the biases of the various sources evaluating minority employees. To understand the difference between $\widehat{\text{BLP}}(X)$ and $\text{BLP}^{*}(X)$, suppose there are additional, unmodeled sources that provide information about majority employees, but the DM doesn’t hold any unconscious bias against them. In the limit, the DM’s model about majority employees is accurate, and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ represents the persistent wage gap between minority and majority employees, holding the profile of evaluations fixed.
3.2 General case
This section characterizes the distortionary effects of misspecified learning in general environments.
To begin, write
for the covariance matrix between familiar sources and its counterpart between familiar and unfamiliar sources, respectively. Denote the vector $[\Delta_{1},\cdots,\Delta_{M}]^{\top}$ of misspecifications associated with familiar sources by $\widetilde{\Delta}$, and the vector $[\Delta_{M+1},\cdots,\Delta_{N}]^{\top}$ of misperceptions about unfamiliar sources by $\widehat{\Delta}$.
@@ 3.2 General case
For any symmetric and positive definite matrix $\Sigma^{*}\in\mathbb{R}^{N\times N}$, $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ and $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$.
Theorem 1 prescribes a two-step procedure for aggregating the misspecifications associated with familiar sources into distortions of the DM’s long-run beliefs and prediction. The first step weighs the misspecifications associated with familiar sources ($\widetilde{\Delta}$) by their covariances ($A$). The second step converts these weighted sums into the DM’s misperceptions about unfamiliar sources and the distortion in his prediction of the state, using their respective covariances between familiar sources ($C$ and $\mathbf{1}_{M}$). Notably, the distortion of the prediction depends only on misspecified sources, but not sources with initially unfamiliar biases or the opportunity to learn about their biases over time, a property we call DOOM. Misperceptions about unfamiliar sources satisfy a weaker form of DOOM, or WDOOM, in that they depend on the covariances between these sources and familiar ones, but not on the covariances between unfamiliar sources. We examine the implications of (W)DOOM in Section 4.
In Section 1, we discussed the potential challenges associated with studying multi-sourced misspecified learning with general covariance structures. Given this, the results demonstrated by Theorem 1 are notably sharp. What is more remarkable is the simplicity of their proof, which exploits the fundamental connection between learning source biases and making corresponding predictions. We present the proof below.
Proof of Theorem 1.
Our DM moves in two stages. In the first stage, he forms beliefs about the biases of unfamiliar sources, which amounts to solving the KL minimization problem. In the current linear-Gaussian environment with a known covariance matrix, the KL minimizer is nothing but the constrained least squares estimator of source biases (at population level). [^7] To establish the connection formally, consider, for each $t$, the vector expression $X=b^{*}+(\omega\mathbf{1}_{N}+\epsilon).$ In Section 3.1 we have defined the variance matrix $\Sigma^{*}:=Var[\omega\mathbf{1}_{N}+\epsilon]$. Since $\Sigma^{*}$ is positive definite, define $\epsilon_{u}:={\Sigma^{*}}^{-1/2}[\omega\mathbf{1}_{N}+\epsilon]$, and note that $Var[\epsilon_{u}]\equiv\mathbb{E}[\epsilon_{u}\epsilon_{u}^{\top}]=Identity.$ So the above vector expression becomes ${\Sigma^{*}}^{-1/2}X_{t}={\Sigma^{*}}^{-1/2}b^{*}+\epsilon_{ut},$ where $t$ indices the $t$th observation. Define the constrained least-square estimator for $b^{*}$ as $\mbox{argmin}_{b\in\Theta}\frac{1}{T}\sum_{t=1}^{T}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right)^{\top}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right).$ The objective function is equivalent to $\frac{1}{T}\sum_{t=1}^{T}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right)^{\top}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right)$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\left({\Sigma^{*}}^{-1/2}b^{*}+\epsilon_{ut}-{\Sigma^{*}}^{-1/2}b\right)^{\top}\left({\Sigma^{*}}^{-1/2}b^{*}+\epsilon_{ut}-{\Sigma^{*}}^{-1/2}b\right)$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\left[\left(b^{*}-b\right)^{\top}{\Sigma^{*}}^{-1}\left(b^{*}-b\right)+2\left(b^{*}-b\right)^{\top}{\Sigma^{*}}^{-1/2}\epsilon_{ut}+\epsilon_{ut}^{\top}\epsilon_{ut}\right].$ To solve the minimization problem at the population level, either take $T\to\infty$, or take the expectation at $T=1$. Either way the result is the same, that the population level objective function is just $\left(b^{*}-b\right)^{\top}{\Sigma^{*}}^{-1}\left(b^{*}-b\right).$ Minimizing the population level objective function is simply (2). To solve, we rewrite the KL minimization problem as
@@
$\sum_{i=1}^{N}\Sigma^{*-1}_{ji}\Delta_{i}=0.$
Write the covariance matrix $\Sigma^{*}$ as $\footnotesize\left[\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right]$, where $A$ and $C$ are defined above. Applying the block matrix inverse formula to $\Sigma^{*}$ yields
hence the system of first-order conditions can be rewritten as
Multiplying both sides of the last equation by $D-CA^{-1}B$, we obtain $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ as the DM’s misperceptions about the biases of unfamiliar sources.
In the second stage, the DM uses the result obtained in the first stage to form the best linear predictor of the true state. Since the former is the best constrained linear estimator of source biases, distortions thereof — which have already been optimized in the first stage — should not further distort the prediction in the second stage. Formally, by substituting $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ into the block covariance matrix and simplifying, we obtain
$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}+A^{-1}B(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\\ \hline\cr-(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\end{array}\right]$ (9)
Consequently,
$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta},$
where the last term depends only on misspecified sources and so satisfies DOOM. ∎
Two remarks are in order. First, we assume unit covariance between the true state and the signals to simplify exposition. In general, the distortion of BLP equals $\alpha_{M}^{\top}A^{-1}\widetilde{\Delta}$, where $\alpha_{M}$ is the vector of covariances between misspecified sources and the true state.
Second, in the above derivation, the linearity of both the first-stage estimator and the second-stage predictor is key, while the normality assumption is only used to establish the equivalence between minimizing KL and constrained least squares. In Section 4, we explore the consequence of this observation for alternative decision situations.
3.3 The case of independent errors
This section delves deeper into the roles of familiar sources in the formation of distortions by examining a simple environment in which the errors associated with the signals are jointly independent. Many of our key economic and methodological insights can be effectively illustrated within this simple framework.
To begin, write $\nu_{i}$ for the signal-to-noise ratio $1/v_{i}$, or absolute precision, of source $i$. For each $i\in\mathcal{M}$, define
$\gamma_{i}\coloneqq\frac{\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$
as the relative precision of source $i$ compared to other misspecified sources. The term
$\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\gamma_{i}\Delta_{i}$
captures the total misspecification across the sources in $\mathcal{M}$, weighted by their relative precision.
The next proposition shows that $\overline{\Delta}$ fully captures the distortionary effects of misspecified learning in the current context.
Proposition 1.
When errors are jointly independent:
  1. (i)
    The DM’s misperception about the bias of any unfamiliar source $i\in\mathcal{S}-\mathcal{M}$ is given by $\Delta_{i}=\overline{\Delta}$, and the distortion of his long-run prediction is given by $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
  2. (ii)
    For each $i\in\mathcal{M}$, $\partial\overline{\Delta}/\partial\Delta_{i}=\gamma_{i}>0$ and $\partial\overline{\Delta}/\partial\nu_{i}=(\Delta_{i}-\overline{\Delta})/(1+\sum_{j\in\mathcal{M}}\nu_{j})$, which is positive if and only if $\Delta_{i}>\overline{\Delta}$.
To develop intuitions, suppose that the manager in our leading example begins with an erroneous belief that two familiar sources are both impartial, although in reality one is biased against minorities and the other is in favor of them, that is, $b_{1}^{*}<0<b_{2}^{*}$ and $\tilde{b}_{1}=\tilde{b}_{2}=0$. Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences. There is a third, unfamiliar, source, which is impartial in reality. To gauge its position, the manager uses familiar sources as yardsticks. If source one is used as the yardstick, the third source appears favorable to minorities. If, instead, source two is used as the yardstick, then the third source seems hostile. One potential approach is to find a middle ground, that is, a position between $b_{1}^{*}$ and $b_{2}^{*}$, and use it as a neutral point to gauge the position of the unfamiliar source. The more biased source $i$ is than source $j$ or the more precise it becomes relative to source $j$, the closer the neutral point is to $b_{i}^{*}$. These are the exact properties satisfied by our distortionary metric $\overline{\Delta}$.
In the case where $\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}<0$, the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities of magnitude $\overline{\Delta}=|\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}|$, and it is mirrored in his beliefs about unfamiliar sources through misspecified learning. In the long run, the DM underestimates the bias of any unfamiliar source that is adversarial toward minorities while overestimating the bias of any unfamiliar source that is favorable to them, assuming that $\overline{\Delta}$ is not excessively large. He also underpredicts minority employees’ productivity and pays them less than under the correctly specified model by $\overline{\Delta}$.
Part (ii) of Proposition 1 explores the comparative statics of $\overline{\Delta}$. The result $\partial\overline{\Delta}/\partial\Delta_{i}>0$ $\forall i\in\mathcal{M}$ is the most intuitive. It simply says that correcting the unconscious bias associated with any familiar source (bringing $\Delta_{i}$ closer to zero) reduces its impact on the overall distortionary effect. This change mitigates the overall distortion if $\overline{\Delta}$ and $\Delta_{i}$ have the same sign; otherwise it exacerbates the overall distortion.
The comparative statics regarding source precision are more nuanced. As we increase the absolute precision of one source, we also increase its relative precision while decreasing the relative precision of the other source. As a consequence, the DM takes the first source more seriously in the learning process, and misspecifications of its bias distort the DM’s long-run beliefs and behavior more heavily than before. The opposite happens to the second source. Depending on the exact nature of misspecifications, the outcome is either an exacerbation of the distortionary effect, or a mitigation thereof. In the example described above, $\Delta_{1}>\overline{\Delta}>0>\Delta_{2}$. Consequently, raising $\nu_{1}$ by a small amount exacerbates the distortionary effect, while raising $\nu_{2}$ by a small amount does the opposite. Note that this ambiguity hinges on the multiplicity of misspecified sources. If, instead, only source one is misspecified, then $\Delta_{1}>\overline{\Delta}=\nu_{1}\Delta_{1}/(1+\nu_{1})$. Raising $\nu_{1}$ by a small amount unambiguously exacerbates the distortion of the DM’s long-run beliefs and prediction in this case.
Proof of Proposition 1.
The proof of Proposition 1, like many to come, exploits a simple observation: in our context, the covariance matrix $\Sigma^{*}$ of the signals is the sum of the covariance matrix $\Omega^{*}$ of the errors and a rank-one update $\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ that reflects the common state of the world. This model feature allows us to utilize matrix algebra results concerning rank-one updates of matrices — such as the Sherman-Morrison formula — to characterize the long-run outcome.
Fact 1 (Sherman and Morrison 1950).